# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v1

@@ 2.1.2 Methods
Paper Main contribution Arch. Exit policy Application Section
Panda et al. (2016) Preliminary Early Exit work CNN Confidence Image classification 2.1.1
Teerapittayanon et al. (2016) First end-to-end network CNN Entropy Image classification 2.1.1
Bolukbasi et al. (2017) Combination with nets ensemble CNN Learned policy Image classification 2.1.1
Huang et al. (2018) Multi-scale architecture CNN Confidence + budget cost Image classification 2.1.1
Kaya et al. (2019) Slight variation wrt BranchyNet CNN Confidence Image classification 2.1.1
Li et al. (2019) Rescale gradient magnitude CNN Confidence Image classification 2.1.2
Passalis et al. (2019) Bag-of-features + single classifier CNN - Image classification 2.1.2
Phuong and Lampert (2019) Introduce self-distillation CNN - Image classification 2.1.2
Wang et al. (2019) Dynamic loss-weight modification CNN Learned policy Image classification 2.1.2
Wang et al. (2019) Frame dropping according to budget CNN Custom optimiz. problem Video analysis 2.1.3
Zhang et al. (2019) Improves self-distillation loss CNN - Image classification 2.1.2
Chen et al. (2020) Exit problem seen as variational Bayes CNN Learned policy Image classification, few shot classification, image denoising 2.1.2
Fang et al. (2020) Applies Early Exits to video analysis CNN Entropy Activity recognition, video understanding, traffic surveillance 2.1.3
Laskaridis et al. (2020) Split computations between on-device and cloud CNN Confidence Image classification 2.1.3
Laskaridis et al. (2020) First to investigate optimal exit positioning CNN Confidence, entropy Image classification 2.1.2
Passalis et al. (2020) Bag-of-features + single classifier CNN - Image classification 2.1.2
Scardapane et al. (2020) Weighting method to estimate exit confidence CNN Learned policy Image classification 2.1.2
Xing et al. (2020) Application on compressed image enhancement CNN quality confidence Quality enhancement 2.1.3
Yang et al. (2020) Processes images at a coarser scale first CNN Confidence Image classification 2.1.1
Bakhtiarnia et al. (2021) Use of curriculum learning to train CNN Entropy Image classification 2.1.2
Ghodrati et al. (2021) Offline frame sampling strategy with Early Exits CNN Learned policy Video analysis 2.1.3
Ju et al. (2021) Early exits for video analytics CNN Learned policy Video analysis 2.1.3
Liu et al. (2021) Early Exits for semantic segmentation CNN Confidence Image segmentation 2.1.3
Bakhtiarnia et al. (2021) Use of curriculum learning to train CNN Entropy Image classification 2.1.2
Ghodrati et al. (2021) Offline frame sampling strategy with Early Exits CNN Learned policy Video analysis 2.1.3
Ju et al. (2021) Early exits for video analytics CNN Learned policy Video analysis 2.1.3
Liu et al. (2021) Early Exits for semantic segmentation CNN Confidence Image segmentation 2.1.3
Passalis and Tefas (2021) Applies BoF Early Exits to face recognition CNN - Face Recognition 2.1.3
Shao et al. (2021) Early exits graph neural network for point cloud GNN - point cloud classification 2.1.3
Tan et al. (2021) Threshold determination as non-convex problem CNN Learned policy Image classification 2.1.2
[…]
@@
Paper Main contribution Arch. Exit policy Application Section
Wang and Li (2021) Dense knowledge distillation for each exit from all the later exits CNN Confidence Image classification 2.1.2
Wang et al. (2021) Image elaboration at different scales with Early Exits Transf. Confidence Image classification 2.1.1
Wołczyk et al. (2021) Reuse the output of internal classifiers as inductive prediction CNN Confidence Image classification 2.1.2
Bakhtiarnia et al. (2022) Early exits audiovisual crowd counting Transf. - Image classification 2.1.3
Han et al. (2022) The loss is provided by a weighting network CNN Confidence Image classification 2.1.2
Kouris et al. (2022) Framework for Early Exit semantic segmentation CNN Confidence Image segmentation 2.1.3
Pomponi et al. (2022) Train by weighting the prediction with a trained confidence score CNN Confidence Image classification 2.1.2
Sabet et al. (2022) Temporal Early Exits for video object detection CNN Entropy Object detection 2.1.3
Sun et al. (2022) Weighted policy to alleviate gradient conflict problems CNN Confidence Image classification 2.1.2
Yu et al. (2022) Proposes a solution to the train-test mismatch problem CNN Confidence Image classification 2.1.2
Zhang et al. (2022) Experiments with various self-distillation techniques CNN Confidence Image classification 2.1.2
Han et al. (2023) Decouples the features extraction and the early classification CNN Confidence Image classification 2.1.1
Li et al. (2023) learned component to decide where to place the classifiers CNN - Image classification 2.1.2
Li et al. (2023) Early exit architecture applies to a Spiking Neural Network Spiking NN Confidence Image classification 2.1.3
Liu and Moon (2023) Balance the loss contribution by weight prediction CNN Confidence Image classification 2.1.2
Polina et al. (2023) Early Exits applied to Generative Adversarial Networks GAN quality threshold Image generation 2.1.3
Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1
Xia and Bouganis (2023) Study of uncertainty estimation when it comes to Early Exit CNN window of confidence Image classification 2.1.2
Xu et al. (2023) Self-distillation to train Early Exit ViT models Transf. confidence threshold Image classification 2.1.2
Xue et al. (2023) ViT which has also the ability to read and store tokens Transf. - Image classification 2.1.1
Bajpai and Hanawal (2024) ViT with Early Exits for image captioning Transf. Learned policy Image captioning 2.1.3
[…]
@@
Paper Main contribution Arch. Exit policy Application Section
Görmez and Koyuncu (2024) Early Exits with threshold tailored to the class CNN & Transformer class confidence Image classification 2.1.3
Kim and Lee (2024) Method based on the Markov decision process for the early-exit decision CNN Learned policy Image classification 2.1.2
Liu et al. (2024) Introduce an adapter for representations in a shared exit space Transf. Confidence Image classification 2.1.2
Meronen et al. (2024) Bayesian method to highlight out of distribution samples CNN - Image classification 2.1.2
Regol et al. (2024) Combines quantization and Early Exit dynamic network Transf. Confidence Image classification 2.1.2
Regol et al. (2024) Presents a loss for both accuracy and inference cost Transf. Confidence Image classification 2.1.2
Valade et al. (2024) The exiting is formalized as a classification with a reject option CNN Confidence Image classification 2.1.2
Yang et al. (2024) Early Exits for object detection CNN Entropy Object detection 2.1.3
Zhang et al. (2024) Self distillation with a dynamic generation of the importance weights CNN Confidence Image classification 2.1.2
Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1
[…]
@@ 2.1.3 Applications
Early Exits networks have been application in numerous computer vision tasks. In the context of object detection Yang et al. (2024) proposes AdaDets, an architecture that speeds up inference through the use of Early Exits. The model formulates the bounding box prediction as a random variable following a certain probability distribution. This allows one to estimate the reliability of both the recognition and localization of a detected object. Furthermore, an entropy score is used for all the detections to estimate the certainty of the network toward the estimation, and make a decision on where to stop the inference.
Liu et al. (2021) proposes an adaptation of the Early Exits framework to semantic segmentation bringing two novel solutions to the network: a redesign of the exiting heads to support early segmentation exits and a confidence adaptivity method. Similarly, Kouris et al. (2022) proposes a framework for adapting semantic segmentation networks to support Early Exiting. The paper introduces a two-stage scheme for training the network, starting with an end-to-end exit-aware pre-training of the backbone thanks to a novel exit dropout loss for the extraction of semantically strong features in the early layers of the network. In the second stage, given the frozen backbone, all candidate Early Exits are further trained through a distillation scheme. More recently, Görmez and Koyuncu (2024) improved the semantic segmentation on top of Liu et al. (2021) architecture by including a threshold that is tailored to the class instead of globally, with the idea that some classes are easier to predict than others in this context, an thus Early Exits can save computations for those classes.
@@
The work of Wang et al. (2019) proposes a scheduling solution for Early Exits models, applied to real-time video analysis. It takes into account the case where a limited time budget is available due to constrained computing resources. The problem is formulated as an optimal scheduling problem, where the objective is to maximize a general overall utility. The scheduler learns to decide which frames to process and which to discard, as well as how much computation budget needs to be assigned to a given frame given that there is an overall deadline to keep the computations near real-time. Trying to solve the same task, FlexDNN (Fang et al., 2020) adopts an architecture search-based scheme for efficient on-device video analytics. The proposed solution is able to find the optimal architecture for each Early Exits branch with a trade-off balance between the Early Exits rate and its computational overhead. Moreover, FlexDNN derives the Early Exits insertion plan to identify the optimal number and locations of Early Exits. Also, Ghodrati et al. (2021) proposes an offline simple sampling strategy to process a video in combination with an Early Exits strategy to enable efficient recognition. Ju et al. (2021) proposes a Dynamic Early Exits method with the aim to find the scheduling of exit points with the best performance both in terms of confidence and low delay. The model targets the application of video analytics. At each exit point, the method estimates whether processing the following exit points can improve the performance. The framework consists of three modules that perform the following: jointly save the current estimations of the optimal utility-latency trade-off for each exit point; the exit oracle observes the current estimation as the context at each exit point, and dynamically decides whether to exit or not; if the exit oracle decides to exit the inference, the exit assessment updates and trains the toolbox, and moves on to process the next frame.
@@
Laskaridis et al. (2020) introduces SPINN, a distributed inference system that spreads the computations between device and cloud with a progressive inference method to deliver refined inference across diverse computational settings. Also in this case, a scheduler is introduced to optimize the exit policy with the network parameters, as well as the network splitting position at run time. The Early Exits mechanism allows the interruption of either the local or remote inference when a confident prediction is reached. A fallback to locally available results is also introduced in order to guarantee the responsiveness of the system even in situations of unavailable or corrupted connectivity. Similarly, Li et al. (2020) proposes a framework to use Early Exits to automatically decide the correct partition between on-device and on-edge with the idea of coordinating all the available resources for real-time inference.
Finally, Spiking Early-Exit Neural Networks (SEENN) (Li et al., 2023) is an Early Exits solution that operates in the context of a Spiking Neural Network and allows an adaptively varying number of timesteps from which to sample during inference. The main idea is to adapt the number of timesteps the network processes according to the classification difficulty, resulting in an Early Exit in the time dimension. Two variants are presented, where the stopping policy is determined with either the use of a confidence score thresholding or a Reinforcement Learning-based exit policy.
@@ 2.2.1 Dynamic architectures
In this paragraph, we present all papers in which the computation path of the network is decided on the fly based on the input. Among the first papers to introduce an input-based model routing solution is Bolukbasi et al. (2017) who propose an in-depth ensemble of networks chained in an acyclic computation graph. The image is supplied to the first block composed of Alexnet (Krizhevsky et al., 2012), after which a decision component evaluates whether the image should be further processed by a ResNet block (He et al., 2015), an Inception block (Szegedy et al., 2015), both, or neither (and thus exiting earlier).
@@
Li et al. (2020) introduces an architecture that decides adaptively the scale at which each image gets processed. A scale transform path is decided on the fly by a differentiable conditional gate for end-to-end optimization. For each mode, multi-path propagation and skip connections are also considered. Furthermore, the paper presents the proficiency of the model in the context of semantic segmentation. A similar approach is also adopted by SegBlocks (Verelst and Tuytelaars, 2023), but the variation of resolution happens within the image. The presented method adjusts dynamically the processing resolution of image regions based on their complexity. The image is split into blocks, and a lightweight policy network processes a low-resolution version of the input to decide which portions of the image should be processed at a high resolution. The policy network is trained using Reinforcement Learning, with a reward taking into account both computational complexity and task accuracy. It is interesting to notice that high and low resolution regions are processed by the same Convolutional filters, which therefore perceive objects at different scales.
@@
Filters and channels. Lin et al. (2017) performs dynamic channel pruning according to the input image and current feature maps adaptively. The model consists of a CNN and a decision network. The pruning is performed layer-by-layer and modeled as a Markov Decision Process using Reinforcement Learning for training. The agent judges the importance of each convolutional kernel and conducts channel-wise pruning. The Markov Chain is conditioned on the previous state consisting of the previous feature maps. Also Gao et al. (2019) proposes to accelerate the Convolution operation by dynamically skipping unimportant input and output channels. The contribution of each channel in the layer is predicted, and subsequently, channels are dynamically amplified or suppressed with a parametric function based on the output of the previous layer. In this case, the dynamic channel pruning component is learned within the network through Stochastic Gradient Descent.
GaterNet (Chen et al., 2019) starts from a similar idea of a backbone network used in combination with a global gating module to generate binary gates for selectively activating filters in the backbone network based on each input. In this case, the novelty is based on a discretization technique called Improved SemHash (Kaiser and Bengio, 2018) to enable differentiable training of input-dependent binary gates such that the backbone and the gating network can be trained jointly via backpropagation. A key difference with other solutions is that the decision on each gate (for each filter) is made based on a shared global view of the current input.
@@
Blocks. SkipNet (Wang et al., 2018) present a revision of the Residual Network (He et al., 2015) which includes a gating mechanism to selectively skip Convolutional blocks based on the activations of the previous layer. The block-skipping problem is formulated as sequential decision-making and solved through a hybrid learning algorithm combining both Supervised Learning and Reinforcement Learning. Each group of layers has a gating module, which performs the binary decision to skip or execute the subsequent layer. The gating modules are trained in two stages: first, a pre-trained version is obtained through the use of a softmax relaxation of the binary skipping decision; since the parameters are not directly optimized for discrete selection for inference, a fine-tuning stage is performed using REINFORCE (Williams, 1992) to refine the policy without relaxation. EnergyNet (Wang et al., 2018) adopts a very similar approach with two major variations: the use of a dynamic routing policy based on a hybrid energy loss that captures both computational and data movement costs, and the use of a single Recurrent Neural Network with shared parameters for the skip decision of all the blocks.

### v3

@@ 2.1.2 Methods
Paper Main contribution Arch. Exit policy Application Section Panda et al. (2016) Preliminary Early Exit work CNN Confidence Image classification 2.1.1 Teerapittayanon et al. (2016) First end-to-end network CNN Entropy Image classification 2.1.1 Bolukbasi et al. (2017) Combination with nets ensemble CNN Learned policy Image classification 2.1.1 Huang et al. (2018) Multi-scale architecture CNN Confidence + budget cost Image classification 2.1.1 Kaya et al. (2019) Slight variation wrt BranchyNet CNN Confidence Image classification 2.1.1 Li et al. (2019) Rescale gradient magnitude CNN Confidence Image classification 2.1.2 Passalis et al. (2019) Bag-of-features + single classifier CNN - Image classification 2.1.2 Phuong and Lampert (2019) Introduce self-distillation CNN - Image classification 2.1.2 Wang et al. (2019a) Dynamic loss-weight modification CNN Learned policy Image classification 2.1.2 Wang et al. (2019b) Frame dropping according to budget CNN Custom optimiz. problem Video analysis 2.1.3 Zhang et al. (2019) Improves self-distillation loss CNN - Image classification 2.1.2 Chen et al. (2020) Exit problem seen as variational Bayes CNN Learned policy Image classification, few shot classification, image denoising 2.1.2 Fang et al. (2020) Applies Early Exits to video analysis CNN Entropy Activity recognition, video understanding, traffic surveillance 2.1.3 Laskaridis et al. (2020a) Split computations between on-device and cloud CNN Confidence Image classification 2.1.3 Laskaridis et al. (2020b) First to investigate optimal exit positioning CNN Confidence, entropy Image classification 2.1.2 Passalis et al. (2020) Bag-of-features + single classifier CNN - Image classification 2.1.2 Scardapane et al. (2020a) Weighting method to estimate exit confidence CNN Learned policy Image classification 2.1.2 Xing et al. (2020) Application on compressed image enhancement CNN quality confidence Quality enhancement 2.1.3 Yang et al. (2020) Processes images at a coarser scale first CNN Confidence Image classification 2.1.1 Bakhtiarnia et al. (2021) Use of curriculum learning to train CNN Entropy Image classification 2.1.2 Ghodrati et al. (2021) Offline frame sampling strategy with Early Exits CNN Learned policy Video analysis 2.1.3 Ju et al. (2021) Early exits for video analytics CNN Learned policy Video analysis 2.1.3 Liu et al. (2021b) Early Exits for semantic segmentation CNN Confidence Image segmentation 2.1.3 Bakhtiarnia et al. (2021) Use of curriculum learning to train CNN Entropy Image classification 2.1.2 Ghodrati et al. (2021) Offline frame sampling strategy with Early Exits CNN Learned policy Video analysis 2.1.3 Ju et al. (2021) Early exits for video analytics CNN Learned policy Video analysis 2.1.3 Liu et al. (2021b) Early Exits for semantic segmentation CNN Confidence Image segmentation 2.1.3 Passalis and Tefas (2021) Applies BoF Early Exits to face recognition CNN - Face Recognition 2.1.3 Shao et al. (2021) Early exits graph neural network for point cloud GNN - point cloud classification 2.1.3 Tan et al. (2021) Threshold determination as non-convex problem CNN Learned policy Image classification 2.1.2
@@
Paper Main contribution Arch. Exit policy Application Section Wang and Li (2021) Dense knowledge distillation for each exit from all the later exits CNN Confidence Image classification 2.1.2 Wang et al. (2021b) Image elaboration at different scales with Early Exits Transf. Confidence Image classification 2.1.1 Wołczyk et al. (2021) Reuse the output of internal classifiers as inductive prediction CNN Confidence Image classification 2.1.2 Bakhtiarnia et al. (2022) Early exits audiovisual crowd counting Transf. - Image classification 2.1.3 Han et al. (2022) The loss is provided by a weighting network CNN Confidence Image classification 2.1.2 Kouris et al. (2022) Framework for Early Exit semantic segmentation CNN Confidence Image segmentation 2.1.3 Pomponi et al. (2022) Train by weighting the prediction with a trained confidence score CNN Confidence Image classification 2.1.2 Sabet et al. (2022) Temporal Early Exits for video object detection CNN Entropy Object detection 2.1.3 Sun et al. (2022) Weighted policy to alleviate gradient conflict problems CNN Confidence Image classification 2.1.2 Yu et al. (2022) Proposes a solution to the train-test mismatch problem CNN Confidence Image classification 2.1.2 Zhang et al. (2022) Experiments with various self-distillation techniques CNN Confidence Image classification 2.1.2 Han et al. (2023) Decouples the features extraction and the early classification CNN Confidence Image classification 2.1.1 Li et al. (2023a) learned component to decide where to place the classifiers CNN - Image classification 2.1.2 Li et al. (2023b) Early exit architecture applies to a Spiking Neural Network Spiking NN Confidence Image classification 2.1.3 Liu and Moon (2023) Balance the loss contribution by weight prediction CNN Confidence Image classification 2.1.2 Polina et al. (2023) Early Exits applied to Generative Adversarial Networks GAN quality threshold Image generation 2.1.3 Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1 Xia and Bouganis (2023) Study of uncertainty estimation when it comes to Early Exit CNN window of confidence Image classification 2.1.2 Xu et al. (2023) Self-distillation to train Early Exit ViT models Transf. confidence threshold Image classification 2.1.2 Xue et al. (2023) ViT which has also the ability to read and store tokens Transf. - Image classification 2.1.1 Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1 Bajpai and Hanawal (2024) ViT with Early Exits for image captioning Transf. Learned policy Image captioning 2.1.3
@@
Paper Main contribution Arch. Exit policy Application Section Görmez and Koyuncu (2024) Early Exits with threshold tailored to the class CNN & Transformer class confidence Image classification 2.1.3 Kim and Lee (2024) Method based on the Markov decision process for the early-exit decision CNN Learned policy Image classification 2.1.2 Liu et al. (2024c) Introduce an adapter for representations in a shared exit space Transf. Confidence Image classification 2.1.2 Meronen et al. (2024) Bayesian method to highlight out of distribution samples CNN - Image classification 2.1.2 Regol et al. (2024a) Combines quantization and Early Exit dynamic network Transf. Confidence Image classification 2.1.2 Regol et al. (2024b) Presents a loss for both accuracy and inference cost Transf. Confidence Image classification 2.1.2 Valade et al. (2024) The exiting is formalized as a classification with a reject option CNN Confidence Image classification 2.1.2 Yang et al. (2024b) Early Exits for object detection CNN Entropy Object detection 2.1.3 Zhang et al. (2024a) Self distillation with a dynamic generation of the importance weights CNN Confidence Image classification 2.1.2 Addad et al. (2025) Self-attention to improve feature representation and aggregation CNN - Image classification 2.1.1 Bajpai and Hanawal (2025) Reuse of previous exits responses for exit confidence CNN Confidence Image captioning 2.1.2 Kuhse et al. (2025) Early exit into YOLO architecture CNN - Object detection 2.1.3
@@ 2.1.3 Applications
Early Exits networks have been application in numerous computer vision tasks. In the context of object detection Yang et al. (2024b) proposes AdaDets, an architecture that speeds up inference through the use of Early Exits. The model formulates the bounding box prediction as a random variable following a certain probability distribution. This allows one to estimate the reliability of both the recognition and localization of a detected object. Furthermore, an entropy score is used for all the detections to estimate the certainty of the network toward the estimation, and make a decision on where to stop the inference. Recently, Kuhse et al. (2025) discusses a variation of the YOLO architecture to perform object detection with the option of exiting at early stages, in combination with specifically dedicated optimization algorithms.
Liu et al. (2021b) proposes an adaptation of the Early Exits framework to semantic segmentation bringing two novel solutions to the network: a redesign of the exiting heads to support early segmentation exits and a confidence adaptivity method. Similarly, Kouris et al. (2022) proposes a framework for adapting semantic segmentation networks to support Early Exiting. The paper introduces a two-stage scheme for training the network, starting with an end-to-end exit-aware pre-training of the backbone thanks to a novel exit dropout loss for the extraction of semantically strong features in the early layers of the network. In the second stage, given the frozen backbone, all candidate Early Exits are further trained through a distillation scheme. More recently, Görmez and Koyuncu (2024) improved the semantic segmentation on top of Liu et al. (2021b) architecture by including a threshold that is tailored to the class instead of globally, with the idea that some classes are easier to predict than others in this context, an thus Early Exits can save computations for those classes.
@@
The work of Wang et al. (2019b) proposes a scheduling solution for Early Exits models, applied to real-time video analysis. It takes into account the case where a limited time budget is available due to constrained computing resources. The problem is formulated as an optimal scheduling problem, where the objective is to maximize a general overall utility. The scheduler learns to decide which frames to process and which to discard, as well as how much computation budget needs to be assigned to a given frame given that there is an overall deadline to keep the computations near real-time. Trying to solve the same task, FlexDNN (Fang et al., 2020) adopts an architecture search-based scheme for efficient on-device video analytics. The proposed solution is able to find the optimal architecture for each Early Exits branch with a trade-off balance between the Early Exits rate and its computational overhead. Moreover, FlexDNN derives the Early Exits insertion plan to identify the optimal number and locations of Early Exits. Also, Ghodrati et al. (2021) proposes an offline simple sampling strategy to process a video in combination with an Early Exits strategy to enable efficient recognition. Ju et al. (2021) proposes a Dynamic Early Exits method with the aim to find the scheduling of exit points with the best performance both in terms of confidence and low delay. The model targets the application of video analytics. At each exit point, the method estimates whether processing the following exit points can improve the performance. The framework consists of three modules that perform the following: jointly save the current estimations of the optimal utility-latency trade-off for each exit point; the exit oracle observes the current estimation as the context at each exit point, and dynamically decides whether to exit or not; if the exit oracle decides to exit the inference, the exit assessment updates and trains the toolbox, and moves on to process the next frame.
@@
Laskaridis et al. (2020a) introduces SPINN, a distributed inference system that spreads the computations between device and cloud with a progressive inference method to deliver refined inference across diverse computational settings. Also in this case, a scheduler is introduced to optimize the exit policy with the network parameters, as well as the network splitting position at run time. The Early Exits mechanism allows the interruption of either the local or remote inference when a confident prediction is reached. A fallback to locally available results is also introduced in order to guarantee the responsiveness of the system even in situations of unavailable or corrupted connectivity. Similarly, Li et al. (2020a) proposes a framework to use Early Exits to automatically decide the correct partition between on-device and on-edge with the idea of coordinating all the available resources for real-time inference.
Finally, Spiking Early-Exit Neural Networks (SEENN) (Li et al., 2023b) is an Early Exits solution that operates in the context of a Spiking Neural Network and allows an adaptively varying number of timesteps from which to sample during inference. The main idea is to adapt the number of timesteps the network processes according to the classification difficulty, resulting in an Early Exit in the time dimension. Two variants are presented, where the stopping policy is determined with either the use of a confidence score thresholding or a Reinforcement Learning-based exit policy.
@@ 2.2 Computational Routing
Similar to the Early Exits domain, numerical results for the publications in this section are not reported here. Also the field of Dynamic Routing is characterized by significant variations in test architectures and benchmark datasets, which prevents a meaningful direct comparison of performance. As previously discussed, this heterogeneity renders a fair quantitative analysis infeasible without extensive re-implementation and underscores the critical need for a standardized experimental protocol. A broader analysis of these limitations, and future directions to overcome these, are provided in the Discussion (Section 4).
@@ 2.2.1 Dynamic architectures
In this paragraph, we present all papers in which the computation path of the network is decided on the fly based on the input. Among the first papers to introduce an input-based model routing solution is Bolukbasi et al. (2017) who propose an in-depth ensemble of networks chained in an acyclic computation graph. The image is supplied to the first block composed of Alexnet (Krizhevsky et al., 2012), after which a decision component evaluates whether the image should be further processed by a ResNet block (He et al., 2016), an Inception block (Szegedy et al., 2015), both, or neither (and thus exiting earlier).
@@
Li et al. (2020b) introduces an architecture that decides adaptively the scale at which each image gets processed. A scale transform path is decided on the fly by a differentiable conditional gate for end-to-end optimization. For each mode, multi-path propagation and skip connections are also considered. Furthermore, the paper presents the proficiency of the model in the context of semantic segmentation. A similar approach is also adopted by SegBlocks (Verelst and Tuytelaars, 2023), but the variation of resolution happens within the image. The presented method adjusts dynamically the processing resolution of image regions based on their complexity. The image is split into blocks, and a lightweight policy network processes a low-resolution version of the input to decide which portions of the image should be processed at a high resolution. The policy network is trained using Reinforcement Learning, with a reward taking into account both computational complexity and task accuracy. It is interesting to notice that high and low resolution regions are processed by the same Convolutional filters, which therefore perceive objects at different scales.
@@
Filters and channels. Lin et al. (2017) performs dynamic channel pruning according to the input image and current feature maps adaptively. The model consists of a CNN and a decision network. The pruning is performed layer-by-layer and modeled as a Markov Decision Process using Reinforcement Learning for training. The agent judges the importance of each convolutional kernel and conducts channel-wise pruning. The Markov Chain is conditioned on the previous state consisting of the previous feature maps. Also Gao et al. proposes to accelerate the Convolution operation by dynamically skipping unimportant input and output channels. The contribution of each channel in the layer is predicted, and subsequently, channels are dynamically amplified or suppressed with a parametric function based on the output of the previous layer. In this case, the dynamic channel pruning component is learned within the network through Stochastic Gradient Descent.
GaterNet (Chen et al., 2019b) starts from a similar idea of a backbone network used in combination with a global gating module to generate binary gates for selectively activating filters in the backbone network based on each input. In this case, the novelty is based on a discretization technique called Improved SemHash (Kaiser and Bengio, 2018) to enable differentiable training of input-dependent binary gates such that the backbone and the gating network can be trained jointly via backpropagation. A key difference with other solutions is that the decision on each gate (for each filter) is made based on a shared global view of the current input.
@@
More recently, Wang et al. (2025) presents a module that treats the height and width of the convolutional kernel as learnable parameters, allowing the shape of the kernel to adjust dynamically based on the size of different objects.
@@
Blocks. SkipNet (Wang et al., 2018a) present a revision of the Residual Network (He et al., 2016) which includes a gating mechanism to selectively skip Convolutional blocks based on the activations of the previous layer. The block-skipping problem is formulated as sequential decision-making and solved through a hybrid learning algorithm combining both Supervised Learning and Reinforcement Learning. Each group of layers has a gating module, which performs the binary decision to skip or execute the subsequent layer. The gating modules are trained in two stages: first, a pre-trained version is obtained through the use of a softmax relaxation of the binary skipping decision; since the parameters are not directly optimized for discrete selection for inference, a fine-tuning stage is performed using REINFORCE (Williams, 1992) to refine the policy without relaxation. EnergyNet (Wang et al., 2018b) adopts a very similar approach with two major variations: the use of a dynamic routing policy based on a hybrid energy loss that captures both computational and data movement costs, and the use of a single Recurrent Neural Network with shared parameters for the skip decision of all the blocks.